MCE Craig, Inc., has provided the following information for one of its products for each hour of production: Actual velocity: 100 units (per hour) Move time: 18 minutes Inspection time: 13 minu...

MCE

Craig, Inc., has provided the following information for one of its products for each hour of production:

Actual velocity: 100 units (per hour)
Move time: 18 minutes
Inspection time: 13 minutes
Rework time: 15 minutes

Required:

1. Calculate MCE. If required, round your answer to two decimal places.

2. What is the theoretical cycle time? Calculate MCE using actual and theoretical cycle times. If required, round your answers to two decimal places.

Theoretical cycle time minutes
Actual cycle time minutes
MCE

3. What if waste is reduced by 30 percent?
New waste =  minutes

What is the new MCE? If required, round your answer to two decimal places.

What is the new cycle time?
minutes
Answer #1

ANSWER

1.

Process time          = Total time – Move time – Inspection time – Rework time

Process time = 60 minutes – 18 minutes – 13 minutes – 15 minutes = 14 minutes

MCE = Process time /(Process time + Move time + Inspection time + Rework time)

MCE = 14/(14+18+13+15) = 0.233

2.

Theoretical cycle time = Process time / Number of units in an hour = 14 minutes/100 units = 0.14 minute

Actual cycle time = Total time/Number of units = 60 minutes/100 units = 0.60 minutes

MCE = Theoretical cycle time/Actual cycle time = 0.14/0.60 = 0.233

3.    

New waste = (1-0.30)(18 minutes + 13 minutes + 15 minutes) = 32.2 minutes

       

New MCE = 14/(14 + 32.2) = 0.30

New cycle time = (14+32.2)/100 = 0.462 minutes

----------------------------------------------------------------------------------------------------

***************THANK YOU*****************